1.The intensity of sound is found to be 7.3×10-5 Wm-2 at3.5m.Find the soind intensity if the observer is:

a. 2m

b. 5m

c. 10m


2.Solve for the power in problem number 1 (a,b,c)


1
Expert's answer
2022-02-20T15:45:49-0500
"I=I_0\frac{r^2}{r_0^2}"

(a)

"I=7.3×10^{-5}\:\rm Wm^{-2}*\frac{2^2}{3.5^2}=2.4*10^{-5}\: \:\rm Wm^{-2}"

(b)

"I=7.3×10^{-5}\:\rm Wm^{-2}*\frac{5^2}{3.5^2}=1.5*10^{-4}\: \:\rm Wm^{-2}"

(c)

"I=7.3×10^{-5}\:\rm Wm^{-2}*\frac{10^2}{3.5^2}=6.0*10^{-4}\: \:\rm Wm^{-2}"